What is the mass of a cubic meter of air at 20oC at sea level?

1
Expert's answer
2021-12-17T18:16:19-0500

The mass of a cubic meter (density) of air is given as follows (see https://en.wikipedia.org/wiki/Density_of_air#Dry_air):


"\rho =\dfrac{p}{R^*T}"

where "p = 1.01\times 10^5Pa" is the pressure at sea level (average), "T = (273 + 20\degree C) = 293K" is the temperature, and "R^* = 278J\/kg\/K" is the specific gas constant for air. Thus, have:


"\rho = \dfrac{1.01\times 10^5}{278\cdot 293} \approx 1.24kg/m^3"

Answer. 1.24 kg.